# QuLTSF: Long-Term Time Series Forecasting with Quantum Machine Learning

### v1

@@ Abstract
Long-term time series forecasting (LTSF) involves predicting a large number of future values of a time series based on the past values and is an essential task in a wide range of domains including weather forecasting, stock market analysis, disease outbreak prediction. Over the decades LTSF algorithms have transitioned from statistical models to deep learning models like transformer models. Despite the complex architecture of transformer based LTSF models ‘Are Transformers Effective for Time Series Forecasting? (Zeng et al., 2023)’ showed that simple linear models can outperform the state-of-the-art transformer based LTSF models. Recently, quantum machine learning (QML) is evolving as a domain to enhance the capabilities of classical machine learning models. In this paper we initiate the application of QML to LTSF problems by proposing QuLTSF, a simple hybrid QML model for multivariate LTSF. Through extensive experiments on a widely used weather dataset we show the advantages of QuLTSF over the state-of-the-art classical linear models, in terms of reduced mean squared error and mean absolute error.
@@ 1 INTRODUCTION
Quantum machine learning (QML) is an emerging field that combines quantum computing and machine learning to enhance tasks like classification, regression, generative modeling using the currently available noisy intermediate-scale quantum (NISQ) computers [Preskill, 2018, Schuld and Petruccione, 2021, Simeone, 2022]. Hybrid models containing classical neural networks and variational quantum circuits (VQC’s) are increasingly becoming popular for various machine learning tasks, thanks to rapidly evolving software tools [Bergholm et al., 2018]. The existing QML models for TSF focus on RNN’s [Emmanoulopoulos and Dimoska, 2022, Ceschini et al., 2022]. In time series analysis, recurrent quantum circuits have demonstrated provable computational and memory advantage during inference, however learning such models remains challenging at scale [Binder et al., 2018]. To the best of our knowledge there is no QML model for LTSF problems.
@@
Organization of the paper: Section $2$ provides background information on quantum computing and quantum machine learning. Section $3$ discusses the problem formulation and related work. In Section $4$ we introduce our QuLTSF model. Experimental details and results are given in Section $5$ and Section $6$ concludes the paper.
@@ 2.2 Quantum Machine Learning
The most common QML paradigm refers to a two step methodology consisting of a variational quantum circuit (VQC) or ansatz and a classical optimizer, where VQC is composed of parametrized quantum gates and fixed entangling gates. The classical data is first encoded into a quantum state, using a suitable data embedding procedure. Then, the VQC applies a parametrized unitary operation which can be controlled by altering its parameters. The output of the quantum circuit is given by measuring the qubits. The parameters of the VQC are optimized using classical optimization tools to minimize a predefined loss function. Often VQC’s are paired with classical neural networks, creating hybrid QML models. Several packages, for instance [Bergholm et al., 2018], provide software tools to efficiently compute gradients for these hybrid QML models.
@@ 5.4 Hyperparameters
The number of qubits $N=10$ and number of VQC layers $K=3$ in the hidden quantum layer. Adam optimizer [Kingma and Ba, 2017] is used to train the model in order to minimize the MSE over the training set. Other hyperparameters are adapted from [Zeng et al., 2023]. For more hyperparameters refer to our code.
@@ 5.5 Results
To further validate QuLTSF against the baseline linear models we conduct experiments with varying sequence lengths $L\in\{48,96,192,336,504,720\}$ and plot the MSE results for a fixed smaller prediction length $T=96$ in Fig. 2, and for a fixed larger prediction length $T=720$ in Fig. 3. In all cases QuLTSF outperforms all the baseline linear models.
@@ 5.6 Discussion and Future Work
QuLTSF uses generic hardware-efficient ansatz. Similar to classical machine learning in QML we need to choose the ansatz, if possible, based on dataset and domain expertise. Searching for optimal ansatz is a research direction by itself [Du et al., 2022]. Finding better ansatzes for QML based LTSF models for different datasets is an open problem. One potential way is to use parameterized two qubit rotation gates [You et al., 2021].
Other possible future direction is to use efficient data preprocessing, for example reverse instance normalization [Kim et al., 2021] to mitigate the distribution shift between training and testing data. This is already being used in the state-of-the-art transformer based LTSF models like PatchTST [Nie et al., 2023] and MTST [Zhang et al., 2024]. These models also show better performance than the linear models in [Zeng et al., 2023]. Interestingly, our simple QuLTSF model outperforms or comparable to these models in limited settings. For instance, for the setting $(L=336,T=720)$ MSE of PatchTST, MTST and QuLTSF are $0.320$, $\underline{0.319}$ and $\mathbf{0.315}$ respectively. For the setting $(L=336,T=336)$ MSE of PatchTST, MTST and QuLTSF are $0.249$, $\mathbf{0.246}$ and $\underline{0.248}$ respectively (see Table $2$ in [Zhang et al., 2024] for PatchTST and MTST; and Table 1 for QuLTSF). Thus, QML models with efficient data preprocessing may lead to improved results.
@@ 6 CONCLUSIONS
We proposed QuLTSF, a simple hybrid QML model for LTSF problems. QuLTSF combines the power of VQC’s with classical linear neural networks to form an efficient LTSF model. While the simple linear models outperform the complex transformer based LTSF models, the introduction of hidden quantum layer further improved the performance. This opens up a new direction of applying hybrid QML models for future LTSF research.
@@ ACKNOWLEDGMENTS
This research is sponsored by the Quantum Engineering Programme (QEP), project ID : NRF2021-QEP2-02-P06.

### v2

@@ Abstract
Long-term time series forecasting (LTSF) involves predicting a large number of future values of a time series based on the past values. This is an essential task in a wide range of domains including weather forecasting, stock market analysis and disease outbreak prediction. Over the decades LTSF algorithms have transitioned from statistical models to deep learning models like transformer models. Despite the complex architecture of transformer based LTSF models ‘Are Transformers Effective for Time Series Forecasting? (Zeng et al., 2023)’ showed that simple linear models can outperform the state-of-the-art transformer based LTSF models. Recently, quantum machine learning (QML) is evolving as a domain to enhance the capabilities of classical machine learning models. In this paper we initiate the application of QML to LTSF problems by proposing QuLTSF, a simple hybrid QML model for multivariate LTSF. Through extensive experiments on a widely used weather dataset we show the advantages of QuLTSF over the state-of-the-art classical linear models, in terms of reduced mean squared error and mean absolute error.
@@ 1 INTRODUCTION
Quantum machine learning (QML) is an emerging field that combines quantum computing and machine learning to enhance tasks like classification, regression, generative modeling etc., using the currently available noisy intermediate-scale quantum (NISQ) computers [Preskill, 2018, Schuld and Petruccione, 2021, Simeone, 2022]. Hybrid models containing classical neural networks and variational quantum circuits (VQC’s) are increasingly becoming popular for various machine learning tasks, thanks to rapidly evolving software tools [Bergholm et al., 2018, Broughton et al., 2020]. The existing QML models for TSF focus on RNN’s [Emmanoulopoulos and Dimoska, 2022, Ceschini et al., 2022]. In time series analysis, recurrent quantum circuits have demonstrated provable computational and memory advantage during inference, however learning such models remains challenging at scale [Binder et al., 2018].
@@
Organization of the paper: Section $2$ provides background information on quantum computing and quantum machine learning. Section $3$ discusses the problem formulation and related work. In Section $4$ we introduce our QuLTSF model. Experimental details, results, discussion and future research directions are given in Section $5$ and Section $6$ concludes the paper.
@@ 2.2 Quantum Machine Learning
The most common QML paradigm refers to a two step methodology consisting of a variational quantum circuit (VQC) or ansatz and a classical optimizer, where VQC is composed of parametrized quantum gates and fixed entangling gates. The classical data is first encoded into a quantum state, using a suitable data embedding procedure like amplitude embedding, angle embedding etc. Then, the VQC applies a parametrized unitary operation which can be controlled by altering its parameters. The output of the quantum circuit is given by measuring the qubits. The parameters of the VQC are optimized using classical optimization tools to minimize a predefined loss function. Often VQC’s are paired with classical neural networks, creating hybrid QML models. Several packages, for instance [Bergholm et al., 2018, Broughton et al., 2020], provide software tools to efficiently compute gradients for these hybrid QML models.
@@ 5.4 Hyperparameters
The number of qubits $N=10$ and number of VQC layers $K=3$ in the hidden quantum layer. Adam optimizer [Kingma and Ba, 2017] is used to train the model in order to minimize the MSE over the training set. Batch size is $16$ and initial learning rate is $0.0001$. For more hyperparameters refer to our code.
@@ 5.5 Results
To further validate QuLTSF against the baseline linear models we conduct experiments for the QuLTSF and classical models with varying sequence lengths $L\in\{48,96,192,336,504,720\}$ and plot the MSE results for a fixed smaller prediction length $T=96$ in Fig. 2, and for a fixed larger prediction length $T=720$ in Fig. 3. In all cases QuLTSF outperforms all the baseline linear models.
@@ 5.6 Discussion and Future Work
QuLTSF uses generic hardware-efficient ansatz. Similar to classical machine learning in QML we need to choose the ansatz, if possible, based on dataset and domain expertise. Searching for optimal ansatz is a research direction by itself [Du et al., 2022]. Finding better ansätze for QML based LTSF models for different datasets is an open problem. One potential way is to use parameterized two qubit rotation gates [You et al., 2021].
Other possible future direction is to use efficient data preprocessing, for example reverse instance normalization [Kim et al., 2021] to mitigate the distribution shift between training and testing data. This is already being used in the state-of-the-art transformer based LTSF models like PatchTST [Nie et al., 2023] and MTST [Zhang et al., 2024]. These models also show better performance than the linear models in [Zeng et al., 2023]. Interestingly, our simple QuLTSF model outperforms or comparable to these models in limited settings. For instance, for the setting $(L=336,T=720)$ MSE of PatchTST, MTST and QuLTSF are $0.320$, $\underline{0.319}$ and $\mathbf{0.315}$ respectively. For the setting $(L=336,T=336)$ MSE of PatchTST, MTST and QuLTSF are $0.249$, $\mathbf{0.246}$ and $\underline{0.248}$ respectively (see Table $2$ in [Zhang et al., 2024] for PatchTST and MTST; and Table 1 for QuLTSF). QML based LTSF models with efficient data preprocessing may lead to improved results.
Implementation of QML based LTSF models on the quantum hardware poses significant challenges. As quantum systems grow, maintaining qubit state coherence and minimizing noise become increasingly difficult, leading to errors that degrade computational accuracy and performance [Lidar and Brun, 2013]. Additionally, the barren plateau phenomenon, where the gradient of the cost function vanishes as system size or circuit depth increases, further complicates optimization of VQC’s [McClean et al., 2018]. Addressing these challenges requires innovations in error correction, problem-specific circuit designs, and alternative optimization strategies, all of which are critical for enabling scalable, noise-resilient, and effective QML based LTSF models.
@@ 6 CONCLUSIONS
We proposed QuLTSF, a simple hybrid QML model for LTSF problems. QuLTSF combines the power of VQC’s with classical linear neural networks to form an efficient LTSF model. Although simple linear models outperform more complex transformer-based LTSF approaches, incorporating a hidden quantum layer yielded additional improvements. This is demonstrated by extensive experiments on a widely used weather dataset showing QuLTSF’s superiority over the state-of-the-art classical linear models. This opens up a new direction of applying hybrid QML models for future LTSF research.
@@ ACKNOWLEDGMENTS
This research is sponsored by the Singapore Quantum Engineering Programme (QEP), project ID : NRF2021-QEP2-02-P06.